Perform the indicated operation and simplify.
x^2+x-12 over x^2-5x+6 - x+3 over x-2

Answer by CubeyThePenguin(3113)

(x^2 + x - 12)/(x^2 - 5x + 6) - (x+3)/(x-2)
= ((x+4)(x-3))/((x-2)(x-3)) - (x+3)/(x-2)
= (x+4)/(x-2) - (x+3)/(x-2)
= 1/(x-2)........(x =/= 3)